OpenStudy (anonymous):

write as a single radical, then simplify: square root, of the square root of 243(x+1)

OpenStudy (anonymous):

\[\sqrt{\sqrt{243(x+1)}}\]

OpenStudy (anonymous):

Weird that that just posted here. Was meant to be on the last question.

OpenStudy (anonymous):

\[\sqrt{\sqrt{243(x+1)}}\]

OpenStudy (anonymous):

nice job!

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

\[\sqrt{\sqrt{b}}=\sqrt[4]{b}\] does that help?

OpenStudy (anonymous):

uhm..not really. i dont know what to do. or where to start

OpenStudy (anonymous):

what i wrote is where to start. put \[\sqrt[4]{243(x+1)}\]

OpenStudy (anonymous):

then you next job is to factor 243 and see if it has any fourth powers in it. if so you can pull them out

OpenStudy (anonymous):

if you are a calculating savant you may recognize \[243=3^5\]but my guess is you do not, so you have to factor. i would cheat

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=factor243 look at that! who knew?

OpenStudy (anonymous):

i love my calculator. lol.. and okay im sorta understanding more now

OpenStudy (anonymous):

Nice work, Satellite.

OpenStudy (anonymous):

awesome.! haha

OpenStudy (anonymous):

Meg, just try to get comfortable with exponent rules.

OpenStudy (anonymous):

(Because roots are really just exponents)

OpenStudy (anonymous):

im just learning the ones past 3... :/ dont really know them

OpenStudy (anonymous):

i am fairly sure that your calculator will not factor, so acquaint yourself with wolfram or else do it by hand

OpenStudy (anonymous):

ohh alright. isee.

OpenStudy (anonymous):

Like for this one, you can rewrite it as (243(x+1)) ^ (1/2) ^(1/2) Then you can know that with exponents, if you raise an exponent to another exponent you multiply them. So this gives: (243(x+1))^(1/4) And with exponents, if it's a numerator it's a power, but in the denominator it's a root. So that makes it: \[\sqrt[4]{243(x+1)}\]

OpenStudy (anonymous):

now you have \[\sqrt[4]{3^5(x+1)}\] and having an exponent in the radicand larger than the index is taboo, so you think \[3^5=3^4\times 3\] and \[\sqrt[4]{3^4}=3\] (you do this part in your head) giving \[3\sqrt[4]{3(x+1)}\] as your answer

OpenStudy (anonymous):

really mean you do that in your head. if you see \[\sqrt[3]{x^7}\] you go right to \[x^2\sqrt[3]{x}\]

OpenStudy (anonymous):

here is the thinking: "x goes in to 7 twice, with a remainder of 1" "x^2 comes outside the radical, x^1 =x stays inside"

OpenStudy (anonymous):

ahhh. i totally get that! thanks so much.!!

OpenStudy (anonymous):

really? great! just to make sure though... what is \[\sqrt{50x^5y^3}\]?

OpenStudy (anonymous):

(hint, the index here is 2, even though it is not written)

OpenStudy (anonymous):

what do u mean index?

OpenStudy (anonymous):

btw i meant here is the thinking: "3 goes in to 7 twice, with a remainder of 1" "x^2 comes outside the radical, x^1 =x stays inside"

OpenStudy (anonymous):

in the expression \[\sqrt[4]{abc}\] 4 is the "index" and abc is the "radicand"

OpenStudy (anonymous):

\[\sqrt[3]{x}\] reads "cubed root of x" index is 3

OpenStudy (anonymous):

alright. my teacher is terrrible at explainng this

OpenStudy (anonymous):

\[\sqrt{x}\] reads "square root of x" index is 2 and by convention we do not write it

OpenStudy (anonymous):

yeah most just say "simplify" which is meaningless unless you know the specific rules for "simplest radical form"

OpenStudy (anonymous):

if you have time read this. the rules are here

OpenStudy (anonymous):

the answer... is it \[2\sqrt{25x^5y^3}\] ??

OpenStudy (anonymous):

no you have a few problems here. i will talk it out. \[50=25\times 2=5^2\times 2\] so a 5 comes out and a 2 stays in 2 (the index) goes in to 5 2 times with a remainder of 1 so x^2 comes out and x stays in 2 goes in to 3 1 time with a remainder of 1, so y^1 = y comes out and y stays in you get \[\sqrt{50x^5y^3}=5x^2y\sqrt{2xy}\]

OpenStudy (anonymous):

if you want we can do another example from your homework

OpenStudy (anonymous):

oh my gosh thanks. and okay.. ill give u one more.. hang on

OpenStudy (anonymous):

\[\sqrt[4]{562^4}\]

OpenStudy (anonymous):

oooh that is like asking who is buried in grants tomb

OpenStudy (anonymous):

what is the fourth root of something raised to the power of 4?

OpenStudy (anonymous):

\[\sqrt[4]{x^4}=?\]

OpenStudy (anonymous):

x...

OpenStudy (anonymous):

ohh so is the answer just 562 ?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

next...

OpenStudy (anonymous):

yay! you are really good at explaining things. im so glad you helped me. (:

OpenStudy (anonymous):

100^-3/2 kinda simple....

OpenStudy (anonymous):

yw. be happy to do another one if you think it would help

OpenStudy (anonymous):

can u help with 100^3/2

OpenStudy (anonymous):

yes. first off do you know what \[x^{\frac{3}{2}}\] means?

OpenStudy (anonymous):

uhmm not exactly.

OpenStudy (anonymous):

the exponent is a fraction. it has a numerator of 3 and a denominator of 2 the numerator is the power, and the denominator is the root

OpenStudy (anonymous):

oh okay....

OpenStudy (anonymous):

so this means a) cube 100, then take the square root b) take the square root of 100, then cube second way is easiest

OpenStudy (anonymous):

what is the square root of 100?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

yes. now what is \[10^3\]?

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

exactly and so \[100^{\frac{3}{2}}=\sqrt{100^3}=10^3=1000\] and you are done

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

okay well i have some others like that but theyre just negative.. is there a difference? on how to do them that is.. like 32^-3/5

OpenStudy (anonymous):

oh yes they are different.

OpenStudy (anonymous):

Negative exponents move to the denominator and become positive. So 2^(-1) is the same as 1/2

OpenStudy (anonymous):

the numerator is the power the denominator is the root and the minus sign means take the reciprocal

OpenStudy (anonymous):

so just switch them around...?

OpenStudy (anonymous):

so if you were going to compute \[32^{-\frac{3}{5}}\] i would first compute \[32^{\frac{3}{5}}\] and the flip the result

OpenStudy (anonymous):

so lets go slow

OpenStudy (anonymous):

ohh okay so same process, just flip the answer.? and oka

OpenStudy (anonymous):

first lets compute \[32^{\frac{3}{5}}\]

OpenStudy (anonymous):

the numerator is 2, so second power. the denominator is 5 so 5th root. i prefer method b) from above. what is the 5th root of 32 (make a guess, it will probably be right)

OpenStudy (anonymous):

2 so i worked it sand got 8 as the answer...

OpenStudy (anonymous):

and *

OpenStudy (anonymous):

great! now we have \[32^{\frac{3}{5}}=2^3=8\] right. so now we take care of the minus sign. take the reciprocal of 8 and you are done

OpenStudy (anonymous):

by which i mean just flip it. answer...

OpenStudy (anonymous):

so its... -1/8 ?

OpenStudy (anonymous):

close but you were thinking too hard. just flip it. don't make it negative

OpenStudy (anonymous):

oh alright. haha. (:

OpenStudy (anonymous):

when you raise to a power if the base is positive you NEVER get a negative answer. your doing great

OpenStudy (anonymous):

got another?

OpenStudy (anonymous):

thank you again. (: and no im all good now that you explained it really good. i may need help again some other day. so i may ask u then but goodnight satellite..!

OpenStudy (anonymous):

g'night